hallm:

solve for t 3t+3/5=6

jasonmitchell:

9/5

hallm:

how did you get that tho

jasonmitchell:

solving it

hallm:

how did you solve it

ThisGirlPretty:

@SmokeyBrown do you think you can explain? >~<

jasonmitchell:

sure i can alright

jasonmitchell:

First Subtract 3/5 from both sides. 3t + 3/5 − 3/5 = 6 − 3/5 3t = 27/5 Then Divide both sides by 3. 3t/ 3 = 27/5/3 t = 9/5 Answer: t = 9/5